40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 11 18:11:46 2015, Koji, Summary, LSC, 3f modulation cancellation 
    Reply  Thu Feb 12 23:47:45 2015, Koji, Update, LSC, 3f modulation cancellation 
       Reply  Sat Feb 14 00:48:13 2015, Koji, Update, LSC, 3f modulation cancellation 
          Reply  Sat Feb 14 19:54:04 2015, Koji, Summary, LSC, 3f modulation cancellation beat_setup1.JPGbeat_setup2.JPGelectrical_setup.pdf
             Reply  Sat Feb 14 20:37:51 2015, Koji, Summary, LSC, 3f modulation cancellation beat_pd_response.pdfbeat_nominal.pdf3f_reduction.pdfbeat_3f_reduced.pdf
                Reply  Sat Feb 14 22:14:02 2015, Koji, Summary, LSC, [HOW TO] 3f modulation cancellation freq_gen_box.JPGdelay_line.JPGcable_spec.pdf
                   Reply  Sun Feb 15 16:20:44 2015, Koji, Summary, LSC, [ELOG LIST] 3f modulation cancellation 
                      Reply  Sun Feb 15 20:55:48 2015, rana, Summary, LSC, [ELOG LIST] 3f modulation cancellation 
                         Reply  Mon Feb 16 00:08:44 2015, Koji, Summary, LSC, [ELOG LIST] 3f modulation cancellation 
                   Reply  Mon Feb 16 01:45:12 2015, Koji, Summary, LSC, modulation depth analysis modulation_nominal.pdfmodulation_3f_reduced.pdf
Message ID: 11028     Entry time: Sat Feb 14 00:48:13 2015     In reply to: 11019     Reply to this: 11029
Author: Koji 
Type: Update 
Category: LSC 
Subject: 3f modulation cancellation 

[SUCCESS] The 3f sideband cancellation seemed worked nicely.

- Beat effeciency improved: ~30% contrast (no need for amplification)

- PLL locked

- 3f modulation sideband was seen

- The attenuation of the 55MHz modulation and the delay time between the modulation source was adjusted to
have maximum reduction of the 3f sidebands as much as allowed in the setup. This adjustment has been done
at the frequency generation box at 1X2 rack.

- The measurement and receipe for the sideband cancellation come later.


- This means that I jiggled the modulation setup at 1X2 rack. Now the modulation setup was reverted to the original,
but just be careful to any change of the sensing behavior.

- The RF analyzer was returned to the control room.

- The HEPA speed was reduced from 100% (during the action on the table) to 40%.

ELOG V3.1.3-